Please help me solve this problem: 19-(4m+6)=2+4m-13.
Found 2 solutions by algebra hello, rothauserc:
Answer by algebra hello(55)

19-(4m 6)=2 4m-13
19-4m-6=2 4m-13
-4m-4m=2-13-19 6
-8m=-24
m=24/8
m=3

Answer by rothauserc(4718)

19 -(4m+6) = 2 +4m -13
:
19 -4m -6 = 2 +4m - 13
:
13 -4m = 4m -11
:
24 -4m = 4m
:
8m = 24
:
m = 3
:
substitute for m in the first equation
:
19 -(4(3) +6) = 2 +4(3) -13
:
19 -(12 +6) = 2 +12 -13
:
19 -18 = 1
:
1 = 1
:
our answer checks